OpenStudy (anonymous):

You live in Pennsylvania and your grandparents live in Ohio. When you are in Pennsylvania, you drive an average rate of 55 mph. When you are in Ohio, you drive an average rate of 65 mph. The entire trip of 175 miles takes 3 hours. a) How long does it take to get from the Pennsylvania-Ohio border to your grandparent’s house? Answer in units of hours Can you walk me through how to do this?

OpenStudy (danjs):

2 equations and 2 variables

OpenStudy (danjs):

55*P + 65*O = 175 (Total Miles) and P + O = 3 (Total Time)

OpenStudy (danjs):

P = time in Penn O = Time in Ohio

OpenStudy (danjs):

Solve by substitution maybe, P = 3 - O The answer to part a) is O = ## hrs

OpenStudy (anonymous):

I don't understand what you just did can you break it down more

OpenStudy (anonymous):

okay I used substitution and I got P=2 hours and O=1

OpenStudy (danjs):

k

OpenStudy (anonymous):

Is it Right?

OpenStudy (danjs):

here

OpenStudy (danjs):

\[Total Miles = \frac{ 55 mi }{ hr }*\frac{ Time Penn }{ 1 } + \frac{ 65 mi }{ hr }*\frac{ Time Ohio }{ 1 }\] and Total Time = Hours Penn + Hours Ohio

OpenStudy (danjs):

Which gives you : 175 = 55*P + 65 *O AND 3 = P + O P=hours in penn O = Hours in Ohio

OpenStudy (danjs):

so O = 1 hr P = 2 hr

OpenStudy (danjs):

The answer to part a, is how much time spent driving in OHIO , O = 1 hour

OpenStudy (anonymous):

thank you So much

OpenStudy (danjs):

yep, not to hard if you read the question once slow, then back again and translate to equations